Investigate how the acceleration of free fall varies around the world. How do these varaitions affect the interpretation of experimental results?
Answer #1

Weight of an object is given by:

W = mg

gravitational force between the objects is:

F = GMm /r^2GMm/r^2 = mg

g = GM/r^2

Using the aobe formula the value of g at normal sa level is g = 9.8 m/s^2 and it varied with the distance of earth.

It varies with the height from the center of the earth.

Following the inverse square law it decreases with increase of altuitude.